# Large Language Model Strategic Reasoning Evaluation through Behavioral Game Theory

### v2

@@ 4.1 Game Library Design
    Signaling games: One player (the sender) conveys a signal $s$ to influence the other player’s (receiver’s) beliefs and subsequent strategy.
In addition, we selected the SW10 game from previous research [70] for its ability to distinguish LLM reasoning levels when compared to human samples. More specific details of the games and corresponding prompts are presented in the Appendix.
4.2 Parameter Estimation
Building on the probabilistic data collected, we estimate the key reasoning parameters using the TQRE model.
Step 1: Setting up the Estimation Equation.
We begin by inputting the observed probabilities and utilities derived from the payoff matrices. For example, in a two-player game where each player $i\in\{1,2\}$ has an action set $A_{i}=\{a_{i1},a_{i2}\}$, let the utilities for player $1$ be defined as $u_{1}(a_{11},a_{21})=U_{11},u_{1}(a_{11},a_{22})=U_{12},\dots$.
The expected utility for player $1$ at reasoning level $k$ when choosing action $a_{11}$, is :
$U_{11}^{(k)}$ $=\sum_{a_{-1}\in A_{-1}}\prod_{h=0}^{k-1}\sigma_{-1}^{(h)}(a_{-1})\cdot u_{1}(a_{11},a_{-1}),$
where:
$\sigma_{-1}^{(h)}(a_{-1})$ $=\frac{\exp\left(\lambda_{h}\cdot U_{-1}^{(h)}(a_{-1})\right)}{\sum_{a\in A_{-1}}\exp\left(\lambda_{h}\cdot U_{-1}^{(h)}(a)\right)}.$
then,
$p_{11}^{(k)}$ $=\frac{\exp\left(\gamma\cdot k\cdot\sum_{a_{-1}\in A_{-1}}\prod_{h=0}^{k-1}\sigma_{-1}^{(h)}(a_{-1})\cdot u_{1}(a_{11},a_{-1})\right)}{\sum_{a_{1j}\in A_{1}}\exp\left(\gamma\cdot k\cdot\sum_{a_{-1}\in A_{-1}}\prod_{h=0}^{k-1}\sigma_{-1}^{(h)}(a_{-1})\cdot u_{1}(a_{1j},a_{-1})\right)}.$
Step 2: Aggregating Probabilities Across All Levels.
The overall probability of selecting $a_{11}$ in the population is computed as:
$p_{11}=\sum_{k=0}^{K_{\text{max}}}f_{k}\cdot p_{11}^{(k)},\text{where }f_{k}=\frac{\tau^{k}e^{-\tau}}{k!}$
Step 3: Maximum Likelihood Estimation (MLE).
Building on the aggregated probabilities derived in Step 2, we now estimate the parameters $\tau$ (mean reasoning level) and $\gamma$ (contextual scaling) using MLE. For each game $g$, the likelihood function is constructed based on the observed strategy frequencies and the predicted probabilities derived from the TQRE model:
$\ln L_{g}(\tau,\gamma)=\sum_{i\in N_{g}}\sum_{j=1}^{J_{g}}c_{ijg}\ln(p_{ijg}(\tau,\gamma)),$
where $c_{ijg}$ is the observed frequency of player $i$ choosing strategy $j$, and $p_{ijg}(\tau,\gamma)$ is the predicted probability from a logit response function based on expected utilities (Step 1). Parameters are estimated separately for each game to capture context-specific reasoning.
4.3 Behavior Evaluation
As the final step of this framework, we draw a comparison between the vanilla model and models enhanced with external mechanisms, such as incorporating demographic features as system context or using CoT prompting to boost reasoning levels. When introducing a boosting mechanism, we aim to evaluate how it influences the models’ ability to make contextually appropriate decisions in strategic scenarios, particularly in games with practical implications.
With the inclusion of demographic features, we apply regression analysis to uncover the relationships between reasoning levels and individual features. These relationships provide insights into whether the model exhibits biases stemming from the training data. Such analyses are crucial for understanding the broader implications of model design and its alignment with fairness and representational objectives.
5 Evaluation and Results
Using the proposed framework, we evaluated 22 SOTA LLMs under three settings: baseline, CoT-embedded, and demographic-feature-embedded, utilizing the game library. The results reveal that while some advanced models, such as DeepSeek-R1, excel in certain tasks, they do not consistently outperform all games with respect to reasoning capability. In some cases, models with modest size and lower capabilities demonstrate comparable or even reasoning levels. Furthermore, our analysis uncovered significant biases in minority demographic features, including religion, marital status, and sexual orientation. The complete results are provided in the Appendix. Our findings also show that the knowledge distillation could lead to further bias as shown in DeepSeek-R1 distilled models.
5.1 Experimental Design
Models: We evaluate reasoning capabilities of general-purpose language models across different scales. The objective of our study is to examine whether larger models or those excelling in individual benchmarks demonstrate superior strategic reasoning. The selected models include GPT-4o-2024-08-06 [49], GPT-o1-preview [50], GPT-o3-mini [63], DeepSeek-R1 [36], DeepSeek-V2.5 [21], DeepSeek-V3 [22], gemma-2-27b-it [72], Gemini-2.0-Flash-Thinking [71], Granite-3.1-8B-Dense [61], Granite-3.1-3B-MoE, Claude-3-Opus [2], internlm2_5-20b-chat [7], Meta-LLaMA-3.1-405B-Instruct [35], Meta-LLaMA-3.1-8B-Instruct, QwQ-32B-Preview [73, 78], and glm-4-9b-chat [32]. In addition, we selected the DeepSeek-R1 Distilled model on LLaMA-70B, LLaMA-8B, Qwen-32B, Qwen-14B, Qwen-7B, and Qwen-1.5B. These models are selected to examine whether distillation introduces additional biases in the inherited model. Selected models specialize in tasks involving reasoning, mathematics, and logic understanding. All models are set with a temperature of 1, top-p of 1, and no repetition penalty, enabling unrestricted generation while ensuring result consistency.
Comparisons: The baseline for all models is performing the games in their vanilla form. Subsequently, all models without internal CoT are augmented with a zero-shot CoT in the prompt. Additionally, demographic features are introduced through prompts as system knowledge across all models, and the same games are performed under these conditions. Finally, we incorporate CoT prompting with embedded demographic features as system knowledge to evaluate their combined effect. Each individual game was tested 30 times to ensure meaningful and non-random results.
Embedded Demographic Features: The demographic features include 10 distinct socio-demographic groups as shown in Table 1. These personas represent a diverse demographic landscape. We generate the distribution by randomly assigning foundational demographic features to LLMs and recording their responses.
@@ 5.1 Experimental Design
Table 1: The Demographic Features across 10 socio-demographic groups.
Group Demographic Features
@@
Education Level below secondary, graduate or above, and in between
@@
5.2 Baseline Evaluation

### v3

@@ 3.1 Game Library Design
    Signaling: The sender sees both real and fake matrices; the receiver must infer the true game. Payoffs reflect sender/receiver misalignment.
Table 2: Representative matrices for incomplete information games
10,10 5,2
7,5 3,3
(a) Bayes type 1
8,8 6,3
5,4 2,2
(b) Bayes type 2
5,5 2,1
3,2 1,0
(c) Signaling (sender)
4,4 6,3
2,3 1,2
(d) Signaling (receiver)
We also include the SW10 matrix from [54], a benchmark for identifying human reasoning levels. The complete payoff matrix tables are attached in the Appendix. While our experiments use abstract normal-form games $(N,\{A_{i}\},\{u_{i}\})$, these payoffs directly mirror real-world interactions. For example, in a recommender system, the e-commerce platform chooses to recommend one of two items, $a_{1}$ or $a_{2}$, and the user choose to purchase ($b_{1}$) or not ($b_{2}$), yielding utilities $u_{1}(a_{i},b_{j})$ for the platform and $u_{2}(a_{i},b_{j})$ for the user. Similarly, in a negotiation, the proposer offers one of two deals, $a_{1}$ or $a_{2}$, and the responder either accepts ($b_{1}$) or rejects ($b_{2}$), capturing the payoffs of agreement. Thus, payoff matrices capture essential strategic structures in various decision-making scenarios.
3.2 Evaluation Procedure
To quantify the average strategic reasoning depth ($\tau$) from observed behaviors, we proceed as follows:
Step 1. Empirical Choice Frequencies. For each LLM and each game, we collect 30 independent trials and record the frequency $c_{ij}$ with which the model chose action $a_{ij}$.
Step 2. Model-Implied Choice Probabilities. Under the bounded‐rationality framework, a level-$k$ agent’s probability of selecting $a_{ij}$ is given by a logit function over its expected utility $U_{ij}^{(k)}$, scaled by $\lambda_{k}=\gamma\cdot k$. Aggregating across reasoning levels $k=0,1,2,\dots$ with Poisson weight $f_{k}(\tau)$ yields:
$p_{ij}(\tau,\gamma)\;=\;\sum_{k=0}f_{k}(\tau)\;\frac{\exp\bigl(\gamma\,k\,U_{ij}^{(k)}\bigr)}{\sum_{a\in A_{i}}\exp\bigl(\gamma\,k\,U_{ia}^{(k)}\bigr)},\quad f_{k}(\tau)=\frac{\tau^{k}e^{-\tau}}{k!}.$
Example: In a two-player game where each player $i\in\{1,2\}$ has actions $A_{i}=\{a_{i1},a_{i2}\}$ and Player 1’s utilities satisfy
$u_{1}(a_{11},a_{21})=U_{11},\quad u_{1}(a_{11},a_{22})=U_{12},\quad u_{1}(a_{12},a_{21})=U_{21},\quad u_{1}(a_{12},a_{22})=U_{22}$
, one computes $U_{11}^{(k)}$ by summing each payoff $U_{1j}$ weighted by the probability that the opponent (a level-$(<k)$ mixture) plays the corresponding action. That expected utility then enters the probability function to yield $p_{11}^{(k)}$, and the mixture over $k$ produces $p_{11}(\tau,\gamma)$.
Step 3. Maximum Likelihood Fit. We estimate $(\tau,\gamma)$ by maximizing the log-likelihood of the observed counts under the model:
$\max_{\tau,\gamma}\;\sum_{i,j}c_{ij}\,\ln p_{ij}(\tau,\gamma)$
, where $c_{ij}$ is the count of times action $a_{ij}$ was chosen. Numerical optimization yields the values of $\tau$ that best fit the LLM’s choice distribution.
3.3 Model Selection.
Given the rapid evolution of LLMs in both open- and closed-source communities, we select representative models from each to demonstrate our framework’s effectiveness and highlight behavioral differences. While including more models is desirable, our selection captures common trends and supports meaningful evaluation. The selected models include GPT-4o [37], GPT-o1-preview [38], GPT-o3-mini [49], DeepSeek-R1 [27], DeepSeek-V2.5 [13], DeepSeek-V3 [14], gemma-2-27b-it [56], Gemini-2.0-Flash-Thinking [55], Granite-3.1-8B-Dense [47], Granite-3.1-3B-MoE, Claude-3-Opus [2], internlm2_5-20b-chat [6], Meta-LLaMA-3.1-405B-Instruct [26], Meta-LLaMA-3.1-8B-Instruct, QwQ-32B-Preview [57, 62], and glm-4-9b-chat [24].
3.4 Fit Diagnostics and Chance Baseline
We evaluate the maximum log-likelihood (MLL) for each model–game pair and contextualize these against a uniform chance baseline. For simultaneous games with $m$ possible actions for Player 1 and $n$ for Player 2, the baseline log-likelihood is defined as
$\mathrm{MLL}_{\text{chance}}=-\ln(mn).$
For example,
$2\times 2$ $\Rightarrow\mathrm{MLL}_{\text{chance}}=-\ln(4)=1.386,$
$3\times 3$ $\Rightarrow\mathrm{MLL}_{\text{chance}}=-\ln(9)=2.197.$
For sequential games, since only the first mover’s action is analyzed, the baseline reduces to
$\mathrm{MLL}_{\text{chance}}=-\ln(m),$
where $m$ is the number of available actions (e.g., $m=3\Rightarrow 1.099$). Values close to this baseline indicate random or level-0–like behavior, whereas higher (less negative) MLL values reflect better fit quality. Appendix B reports per-model fits and baselines, complementing Table 3.
4 Evaluation Results and Comparison
In this section, we present reasoning depth evaluations, followed by an analysis of reasoning variation across game types through models’ reasoning chains. By identifying distinct reasoning styles, we provide both quantitative and qualitative insights into how LLMs reason in interactive settings.
4.1 Strategic Reasoning Capabilities Across Models
Table 3 presents the reasoning depths evaluated through the framework for each model across games. Among all models, GPT-o1, GPT-o3-mini, and DeepSeek-R1 outperform others, with GPT-o1 ranking first in 6 games, GPT-o3-mini leading in 7 games, and DeepSeek-R1 securing the top position in 5 games. Some games have co-leaders, where multiple models achieve the highest evaluated strategic reasoning level. In these cases, certain LLMs consistently select strategies that align with theoretically robust principles while also demonstrating accurate anticipation of the opponent’s behavior. However, not all models achieve this level of capability, indicating variations in reasoning depth. This shows both the strengths of top models in structured settings and the effectiveness of our framework in distinguishing optimal decision-making from suboptimal or inconsistent reasoning. Additionally, our findings reveal significant variations in reasoning capabilities across different game types.
GPT-o1 consistently ranks among the top models in competitive and incomplete-information games, indicating a strong capacity for goal-oriented and adversarial reasoning. DeepSeek-R1 excels in cooperative and mixed-motive settings, likely due to its reinforcement learning-based optimization for social interaction and mutual benefit. GPT-o3-mini performs robustly across all game types, particularly in cooperative, mixed-motive, and incomplete-information scenarios, reflecting a balanced ability to manage trade-offs, adapt to uncertainty, and apply probabilistic inference.
The performance gap between stronger models, such as DeepSeek-R1, GPT-o1 and GPT-o3-mini, and weaker ones, like DeepSeek V2.5, DeepSeek V3, and GPT-4o, is notable across most games. Nevertheless, advanced models do not always dominate all game types. Smaller models like Gemma-2-27B and LLaMA-3.1-8B sometimes match or outperform larger counterparts, and DeepSeek-R1 is a top model despite not being the largest. Additionally, derived from a much smaller base model, R1-distilled models can outperform significantly larger models  [^1] Full results in Appendix B, Table 7 . This suggests that model size is not the sole factor in reasoning quality. Instead, context, model structure, and training data likely play a crucial role, highlighting the strong influence of contextual adaptation on LLMs’ reasoning abilities.
Table 3: Strategic reasoning depth ($\tau$) across different game settings
Complete Information Game Incomplete Information Game
Competitive
Games
Cooperation
Games
Mixed-motive
Games
Seq.
Games
Bayesian
Games
Signaling
Games
BL HS LS BL HP AP BL H-Pun L-Pun p=0.5 p=0.9
Note: Bold values indicate the highest strategic reasoning level ($\tau$) observed for that game. Abbreviations: HS = High Stake, LS = Low Stake, BL = Baseline, HP = High Payoff, AP = Asymmetric Payoff, H-Pun = High Punishment, L-Pun = Low Punishment.
4.2 Empirical Exploration for Variations of Leading LLM in Different Game Types
To investigate why certain advanced models perform well in specific game types while others do not, we focus our analysis on three top-performing models, DeepSeek-R1, GPT-o1, and GPT-o3-mini, in baseline competitive, cooperative, and mixed-motive games.
We gather available internal or summarized reasoning traces from each model and link them with its observed behavioral choices. For each model–game pair, we extract the empirical strategy distribution from 30 independent trials and identify the dominant or mixed-strategy pattern. We then examine the corresponding reasoning traces through qualitative analysis, aligning textual reasoning steps with revealed strategic behavior to evaluate coherence between what the model says and what it does. To ensure robustness and mitigate subjectivity, we collect a diverse set of reasoning traces from different models and anonymize them before analysis. Researchers who coded each CoT trace for reasoning style did so without knowing which model produced it, following a “blind review” procedure that minimizes post-hoc bias. A pre-defined structured coding rubric was used to classify reasoning styles (e.g., maximin, belief-based anticipation, opponent-oriented logic), and cross-validation was performed by having multiple researchers independently code overlapping subsets and compare results for consistency. Finally, we record the completion tokens required by each model to quantify computational demand and analyze whether longer reasoning chains correspond to improved or diminished decision quality.
Based on the summarization in Table 4, we observe that each model’s performance in different game types reflects the strategic logic it tends to apply. While these styles are not mutually exclusive or deterministic, they offer a coherent lens for interpreting variation in performance.
Table 4: LLMs’ reasoning styles and examples
Model Dominant Reasoning Style Representative Excerpts
GPT-o1
Consistent maximin,
Payoff pessimism (risk-averse)
“I select the row that maximizes our minimum payoff.” (competitive game)
“Even though Row 0 gives 8 in the best case, Row 1 guarantees at least 5.” (coordinative game)
"Worst-case payoff for me (assuming the opponent minimizes our payoff): 0" (mixed strategy game)
GPT-o3
Belief-guided decision-making
with fallback to maximin
“Player 2 is likely to choose column 1, so I respond with row 0.” (competitive game)
“Column 2 guarantees the highest minimum payoff… That’s why I selected column 2.”
(mixed motive game)
DeepSeek-R1
Opponent oriented decision,
Lack of adversarial consideration
“Player 2 will likely choose the column that maximizes their own payoff… so I choose row 0.”
(competitive game)
“… neither player benefits from deviating” (coordinative game)
GPT-o1 consistently applies minimax reasoning, evaluating each option by its worst-case outcome. In nearly every reasoning chain we extracted, the model explicitly states, “I will now use minimax to help me make a decision.” This leads to strong performance in competitive games, where minimizing potential losses aligns with an optimal strategy. However, the same strategy becomes overly cautious in cooperative or mixed-motive games, where achieving mutual benefit often requires trust and risk-taking. GPT-o1 also exhibits a highly defensive stance, at times assuming that the opponent’s objective is to minimize its payoff, interpreting interactions as fully adversarial. GPT-o3-mini demonstrates greater flexibility. It primarily engages in belief-based anticipation, attempting to infer the opponent’s likely move and respond accordingly. This allows it to perform well in cooperative and mixed-motive settings, where alignment with the other player is key. Yet under uncertainty, GPT-o3-mini sometimes falls back to the worst-case logic, which can limit its potential in competitive games but also protects against major failures. DeepSeek-R1 adopts a strategy that begins with assumptions about the opponent’s likely action, typically based on the idea that the opponent will pursue their own payoff. This logic works well in cooperative games, where the players’ incentives are aligned, but lacks the adversarial caution required in competitive games. R1 also tends to exhibit strategic trust, assuming that when mutual benefit is unavailable, the opponent will not deviate simply to harm the other player, in contrast to GPT-o1, which at times assumes spiteful intent.
An additional observation relates to token usage, as shown in Appendix B Table 8. Contrary to intuitive expectations, higher token counts in internal reasoning do not correlate with better reasoning. In fact, GPT-o1 and DeepSeek-R1 are leaders in competitive and cooperative games, respectively, while they produce the shortest CoT outputs within their strongest games. We observe that longer reasoning chains often suggest hesitation and uncertainty, rather than deeper insight. Conversely, more concise CoT responses tend to signal higher confidence and more direct strategy choices. Notably, in competitive games, DeepSeek-R1 often exhibits repeated self-doubt in its CoT, but this results in redundant reasoning loops that inflate token usage without improvement. This points to future research on improving token efficiency in CoT generation, such as incorporating RL-based training that encourages decisive and diverse reasoning over repetitive patterns.
4.3 Effect of CoT Prompting
Our empirical study of top-performing models with embedded internal CoT reasoning motivates us to explore whether CoT prompting can also enhance models that do not natively generate internal reasoning. The strategic reasoning level after explicitly adding the CoT prompt is presented in Appendix B Table 9. Intuitively, many believe CoT can likely enhance a model’s reasoning ability. However, our results show that CoT’s impact on strategic reasoning is mixed, with no consistent improvement across all models and game types, which is consistent with prior findings suggesting that the "CoT does not always help" [18, 10], particularly in strategic settings.
For example, Figure 2 compares $\tau$ in the Competitive-Base game with and without CoT prompting. While some models improve, others remain unchanged or even decline. We select two representative cases: with the largest improvement and sharpest decline: Claude-3-Opus and GPT-4o. Table 5 presents excerpts from their CoT reasoning chains to illustrate the underlying dynamics  [^2] Full reasoning chains are in Appendix E.
Figure 2: Comparison between baseline and CoT
Note: Green marks the largest increase; red marks the largest decline
Figure 3: Comparison among human participants and LLMs on the SW game
Table 5: CoT effect on reasoning examples
Model Prompt Excerpt
Example of Improvement:
"I should pick the row that gives me the maximum value among these worst-case…
Wait, I made a mistake… Let me reconsider using the concept of maximin…"
Claude-3-opus
"Row 0 and Row 2 both have a worst-case of –5, but Row 0 offers a potential payoff of 10,
so I choose Row 0."
Example of Decline:
"We assume Player One picks the row that minimizes our payoff, so…"
GPT-4o "Break tie with average… Column 2 is better."
If we look at their gaming strategies, Claude-3-Opus quickly identifies a clear attempt to apply the maximin strategy at an earlier stage, an appropriate approach for the competitive game. It initially makes calculation errors, referring to incorrect cells during its payoff analysis, but the model explicitly identifies and corrects these mistakes mid-chain. This self-correction enables it to recover and ultimately arrive at a robust decision that aligns with its strategic intent. In contrast, GPT-4o starts with a fundamentally flawed assumption, mistakenly believing that the opponent will act adversarially to minimize its payoff. Based on this incorrect premise, the model applies a tie-breaking rule using average values across columns, stuck in a verbose reasoning loop that reinforces the wrong strategy. These again imply that CoT does not guarantee better reasoning. When a model begins with a sound strategy, CoT can support useful self-correction and refinement. However, when the underlying reasoning is misaligned from the beginning, extended chains often lead to repetitive and unproductive loops rather than meaningful improvement, which echoes one of our findings in Section 4.2.
5 Detecting Social Effect with Human
While reasoning depth is useful for evaluating LLMs’ strategic capacity, model sophistication alone is not enough. Real-world deployment requires alignment with contextual goals, ethical constraints, and human norms. To examine these factors, we compare LLM reasoning with human performance in identical games and examine how assigned demographic personas influence model behavior, shedding light on identity-induced strategic shifts and fairness in AI decision-making.
Figure 3 compares all models against human subjects in a single-scenario setting [54, 51], with six models surpassing human performance. However, this ranking should not imply that “a higher score is better”. In many applications, alignment with human reasoning is more critical than outperformance. Our framework supports such evaluation by quantifying reasoning depth across subjects, offering a path to calibrate LLMs for human-aligned behavior. For instance, businesses can use representative human samples to fine-tune models toward user-aligned decisions.
To further evaluate the alignment of LLMs with human strategic reasoning, we construct a diverse set of socio-demographic personas spanning a broad range of individual characteristics, inspired by [29]. These personas are structured into two tiers: foundational demographic features (Panel 1) and advanced identity attributes (Panel 2), as detailed in Table 6.
@@ 5 Detecting Social Effect with Human
Table 6: The Personas across 10 socio-demographic groups that we explore in this study.
Group Persona
Panel 1: Foundational Demographic Features
@@
Education Level
below lower secondary, lower secondary, upper secondary, short-cycle tertiary, bachelor, and graduate degrees
@@
We then perform Ordinary Least Squares (OLS) regressions using the estimated behavioral parameters as dependent variables. The general specification of the regression is:
$\text{Reasoning\_Depth}_{i}=\beta_{0}+\beta_{n}D_{ni}+\epsilon_{i}$
Here, $i$ indexes each observation, $D_{ni}$ is the $n$-th binary demographic indicator, $\beta_{n}$ denotes its estimated effect on the behavioral parameter, and $\epsilon_{i}$ is the error term.
Figure 4: Significant demographic shifts in reasoning depth
Note: Each point reflects the significant change in reasoning depth relative to a reference group through OLS. Reference categories are: Male (gender), Heterosexual (sexual orientation), Age 25–64 (age), Able-bodied (disability status), Single (marital status), Other Religious (religion), and Democrat (political affiliation). Positive (negative) shifts indicate higher (lower) reasoning depth relative to the reference group.
Figure 4 shows estimated shifts in reasoning depth across demographic conditions with statistically significant effects from OLS regressions. A consistent pattern emerges around gender: when prompted with a “female” persona, several large models—such as GPT-4o, InternLM V2, and Claude 3-Opus—exhibit increased reasoning depth, suggesting deeper multi-level reasoning under female identity framing. Sexuality-based variation reveals further heterogeneity. While “homosexual” and “bisexual” personas raise reasoning depth in models like GPT-4o and GLM 4, they lead to significantly reduced $\tau$ in Gemini 2.0 and Granite MoE. These divergent effects highlight both the contextual flexibility of LLM reasoning and the risk of identity-induced attenuation. Other demographic factors can also cause disturbance in reasoning depths. We include full results of this study in Appendix B.
Mechanism note.
Observed persona-induced shifts likely stem from statistical associations in training corpora that link demographic descriptors to behavioral patterns, potentially modulated by reinforcement learning from human feedback (RLHF). Models often deny explicit bias yet reveal implicit adaptation under identical payoffs, indicating latent contextual priors. Our framework surfaces these implicit shifts and quantifies their magnitude for downstream auditing and fairness alignment.
6 Conclusion and Discussion
This study sets a new milestone in evaluating LLM strategic reasoning by assessing while accounting for contexts, addressing the limitations of NE-focused benchmarks. We find that strong performance on standard tasks (e.g., math) does not always guarantee higher reasoning depth. Analysis also reveals the limitations of CoT prompting in strategic settings, where its benefits are not universal and can, at times, impair reasoning. These findings inform ongoing efforts to refine LLM reasoning workflows for complex decision-making. Additionally, our demographic analysis exposes how biases can emerge even in top-performing models, underscoring the need for fairness-aware evaluation.
Effectiveness of CoT. While CoT can enhance the reasoning process in language models, its effectiveness hinges on task comprehension and alignment with intended objectives. Our analysis shows that well-aligned CoT prompts support structured reasoning and improved outcomes, whereas misaligned prompts can amplify errors—even in models trained with RL from Human Feedback (RLHF). DeepSeek-R1 illustrates that, despite strong performance in some games, it underperforms in others, exhibiting repetitive self-questioning and redundant token generation without better solutions. RLHF helps refine strategies, but without the intrinsic ability to evaluate context-dependent strategies, models may remain trapped in ineffective reasoning loops. These findings highlight the need for adaptive, context-aware reasoning and raise a key open question: how can external prompts be better aligned with a model’s internal reasoning processes to ensure robust and interpretable behavior?
Alignment and Fairness in AI Reasoning. Superior strategic reasoning depth does not equate to ethical decision-making. When demographic cues are introduced, models can adjust their strategies based on the given role, indicating that their decisions can encode and present societal biases. This is especially concerning in multi-agent interactions, where solely prioritizing strategic efficiency may lead to inequitable or adversarial outcomes. In interactive settings, LLMs that adjust strategies based on demographic cues risk perpetuating bias in negotiations and resource allocation, leading to real-world inequalities over time. This presents a critical challenge to balance strategic reasoning with fairness. Unconstrained optimization may encourage exploitative tactics, while rigid fairness constraints could hinder adaptability. Future development must move beyond performance metrics to integrate ethical safeguards, ensuring that LLMs engage in fair and justifiable decision-making in dynamic multi-agent environments.
Limitations and Future Directions: While our study sheds light on LLM reasoning patterns and benchmark alignment, the causal link between prompts and decision outcomes remains underexplored. Future work should investigate how individual cognitive skills—such as logic, math, and contextual understanding—interact to shape reasoning capabilities. Moreover, our current framework is semi-dynamic; extending it to real-time multi-agent interactions may reveal new reasoning dynamics and performance disparities. Exploring such settings could provide deeper insights into adaptive LLM behavior and strategy formation.
Broader Impact: This study provides practical insights into how the internal decision-making processes of large language models translate into real-world social outcomes through quantitative evaluations. Unlike explicit discrimination, these biases emerge through adaptive reasoning, making them more difficult to detect and regulate. As organizations increasingly rely on AI to guide complex decisions affecting diverse populations, researchers and practitioners must remain cautious: subtle reasoning biases in sophisticated AI could inadvertently widen existing societal gaps. Addressing these challenges today will ensure that LLMs become inclusive tools promoting societal equity, rather than unintentionally reinforcing historical injustices.
Acknowledgment
We gratefully thank the anonymous reviewers and area chair for their insightful comments and constructive feedback. We also acknowledge the support provided in part by the AMD Center of Excellence at the University of Illinois Urbana–Champaign.
Appendix A Detailed Model Specification
In this appendix we provide a fully detailed statement of the model and then work through a concrete two‐player, two‐action example for clarity.
A.1 Game and Payoff Structure
  • $\mathcal{I}=\{1,\dots,N\}$: set of players.
  • $A_{i}$: finite action set of player $i$.
  • $A_{-i}=\prod_{j\neq i}A_{j}$: joint action space of all players except $i$.
  • $a_{i}\in A_{i}$, $a_{-i}\in A_{-i}$: player $i$’s action and opponents’ profile.
  • $u_{i}(a_{i},a_{-i})$: payoff to $i$ when playing $a_{i}$ against $a_{-i}$.
A.2 Reasoning Levels and Noise
  • Agents draw a reasoning level $k\in\{0,1,2,\dots\}$ from a Poisson distribution with mean $\tau>0$:
    $f_{k}\;=\;\Pr(\text{level}=k)=\frac{\tau^{k}e^{-\tau}}{k!}\,,\quad k=0,1,2,\dots$
  • Each level-$k$ agent has precision (inverse noise)
    $\lambda_{k}=\gamma\,k,\quad\gamma>0.$
    Higher $\lambda_{k}$ means more sharply payoff‐driven (less random) choices.
A.3 Level-$h$ Strategy Distributions
For each opponent $j\neq i$ and each potential level $h$, we assume a level-$h$ strategy
$\sigma_{j}^{(h)}:A_{j}\to[0,1],\quad\sum_{a_{j}\in A_{j}}\sigma_{j}^{(h)}(a_{j})=1.$
This is the probability that a level-$h$ thinker $j$ chooses action $a_{j}$.
A.4 Mixed Marginal Belief
A level-$k$ agent does not know which of the levels $0,\dots,k-1$ its opponents occupy, so it forms a mixture of their level-$h$ strategies. Define the normalized weight
$W_{h}^{(k)}=\frac{f_{h}}{\sum_{\ell=0}^{k-1}f_{\ell}},\quad h=0,\dots,k-1.$
Then the agent’s marginal belief about opponent $j$ is
$\bar{\sigma}_{j}^{(k)}(a_{j})=\sum_{h=0}^{k-1}W_{h}^{(k)}\,\sigma_{j}^{(h)}(a_{j}),\quad a_{j}\in A_{j}.$
A.5 Joint Belief
The notation
$a_{-i}=(a_{1},\dots,a_{i-1},a_{i+1},\dots,a_{N})$
stands for the actions of all players except $i$. For example, if $N=3$ and $i=1$, then $a_{-1}=(a_{2},a_{3})$, meaning “opponent 2 plays $a_{2}$” and “opponent 3 plays $a_{3}$.” Under the assumption that each opponent $j\neq i$ acts independently, in this example, $\Pr(a_{2},a_{3})=\bar{\sigma}_{2}^{(k)}(a_{2})\,\bar{\sigma}_{3}^{(k)}(a_{3})$, In general, the probability of the profile $a_{-i}$ is
$\Pr(a_{-i})\;=\;\prod_{j\neq i}\bar{\sigma}_{j}^{(k)}(a_{j}).$
A.6 Expected Utility at Level $k$
Given belief $\Pr(a_{-i})$, the expected payoff for player $i$ choosing $a_{ij}\in A_{i}$ is
$U_{ij}^{(k)}=\sum_{a_{-i}\in A_{-i}}\Pr(a_{-i})\,u_{i}\bigl(a_{ij},a_{-i}\bigr).$
A.7 Choice Probabilities
Once a level-$k$ agent has computed the expected payoffs $U_{ij}^{(k)}$ for each action $a_{ij}$, it converts these into a probability distribution via a logit rule. The precision parameter $\lambda_{k}$ governs how sharply the agent favors utility-oriented actions:
$p_{ij}^{(k)}=\frac{\exp\!\bigl(\lambda_{k}\,U_{ij}^{(k)}\bigr)}{\sum_{a\in A_{i}}\exp\!\bigl(\lambda_{k}\,U_{ia}^{(k)}\bigr)}.$
When $\lambda_{k}$ is small, payoffs have little influence and choices are nearly uniform; as $\lambda_{k}\to\infty$, the agent approaches deterministic best-response behavior.
A.8 Population-Level Aggregation
Because agents’ reasoning levels $k$ are drawn from the Poisson weights $f_{k}$, the overall probability that a randomly chosen agent plays action $a_{ij}$ is the weighted average of level-specific probabilities:
$p_{ij}=\sum_{k=0}f_{k}\,p_{ij}^{(k)}.$
This mixture captures heterogeneity in both depth of reasoning and choice stochasticity across the population. In implementation, the infinite sum is truncated at a sufficiently large $K$ so that $\sum_{k=0}f_{k}\approx 1$.
Appendix B Additional Results
In this section, we present the complete results for the following model variants across all game settings: vanilla models, vanilla models with CoT prompting, distilled models based on DeepSeek-R1, vanilla models with embedded demographic features, and vanilla models with both CoT prompting and demographic features.
Distilled Model
We evaluate DeepSeek-R1 distilled models across multiple architectures—LLaMA-70B, LLaMA-8B, Qwen-32B, Qwen-14B, Qwen-7B, and Qwen-1.5B—to assess whether the distillation process introduces additional biases in the inherited models. As shown in Table 7, R1-distilled models can outperform significantly larger counterparts, despite originating from smaller base models. This indicates that model size alone does not determine reasoning quality. Instead, factors such as architectural design, training data, and contextual adaptation play a pivotal role in shaping LLM reasoning capabilities.
However, our analysis reveals that knowledge distillation can also introduce new biases, as evidenced in Table 11(b). While distillation enhances reasoning performance—particularly in cooperative and mixed-motive game settings—it also increases sensitivity to demographic features. In the vanilla LLaMA models, even with CoT prompting, bias remains limited, typically affecting only one or two attributes. In contrast, R1-distilled versions of LLaMA-8B and LLaMA-70B exhibit amplified demographic biases. For instance, DSk-R1-LLaMA70B shows marked sensitivity to age and race.
These findings highlight the critical importance of monitoring the distillation process. If the teacher model is trained on malicious or skewed data, these biases may propagate to the student model, potentially resulting in harmful behaviors. More concerningly, even when trained on ostensibly neutral data, hidden biases can be transferred during distillation. This underscores the need for rigorous auditing and bias mitigation when applying knowledge distillation in LLM development.
CoT + Demographic Features
To assess whether explicit step-by-step reasoning affects model fairness, we introduce CoT prompting and analyze its impact on demographic correlations, as reported in Table 12. Overall, we observe that CoT increases the number of demographic variables exhibiting significant influence—even in models that were previously neutral—indicating that CoT can shift how models interpret and weigh user cues. For instance, Figure 5 shows that Gemma-2 exhibits a marked increase in sensitivity to the “Jewish” attribute, while InternLM, initially unbiased, develops religious disparities under CoT. These findings raise important fairness concerns: CoT can both surface latent biases and introduce new correlations, making fairness properties highly sensitive to prompt design. In particular, if a model overemphasizes certain demographic signals, CoT may exacerbate rather than alleviate bias. This highlights the necessity of prompt-specific bias audits and careful monitoring as CoT prompting becomes more widely adopted.